OpenStudy (anonymous):

How do you find sin, cot, and csc of theta if you only have the hypotenuse and adjacent angle?

OpenStudy (amistre64):

either drawing it out or using the pythag trig identities

OpenStudy (amistre64):

cos = adj/hyp sec = 1/cos sin = sqrt(1-cos^2) csc = 1/sin tan = sin/cos cot = 1/tan